For the function y = x2 - 4x - 5
Put the function in the form y = a(x - h)2 + k

Answer by venugopalramana(3286)

For the function y = x2 - 4x - 5
={X^2-2(X)(2)+2^2}-2^2-5
=(X-2)^2-9
A=1....H=2.......K=-9


Put the function in the form y = a(x - h)2 + k